In a group of 1000 male graduate students, how many would you expect to find over 73 inches tall? The mean=70 inches, the standard deviation= 2 inches.

Answer by stanbon(75887)

In a group of 1000 male graduate students, how many would you expect to find over 73 inches tall? The mean=70 inches, the standard deviation= 2 inches.
-----
z(73) = (73-70)/2 = 3/2
---
P(x > 73") = P(z > 1.5) = 0.0660
---
Expected number = 0.0660*1000 = 66
===================================
Cheers,
Stan H.